If 9x^2 - y^2 = 0 then what is the positive value of the x/y ?
Answer by nerdybill(7384)

If 9x^2 - y^2 = 0 then what is the positive value of the x/y ?
.
Notice, I can rewrite:
9x^2 - y^2 = 0
as
(3x)^2 - y^2 = 0
.
This is a "difference of squares", so I can factor as:
(3x-y)(3x+y) = 0
.
The positive value of x/y is then:
3x-y = 0
3x = y
x = y/3
x/y = 1/3 (this is your answer)